OpenStudy (anonymous):

medal to the first person to answer 2(1/8 + 4) - ( 3/16 * 8)

OpenStudy (jojo4eva):

8.25-2.5

OpenStudy (jojo4eva):

5.75

OpenStudy (anonymous):

it has to be in fraction form

OpenStudy (jojo4eva):

ok 5 3/4

OpenStudy (anonymous):

it has to be one of these answers a.11/4 b.65/2 c.21/4 d.27/4

OpenStudy (jojo4eva):

D

OpenStudy (anonymous):

ok can u help me on another question

OpenStudy (jojo4eva):

no probs

OpenStudy (anonymous):

(12-4) +8/2 a.4 b.6 c.8 d.10

OpenStudy (jojo4eva):

You should get C because you subtract 12-4 and u get 8 and then add that to 8 and u get 16 and then you divide 16 by 2 and get 8

OpenStudy (anonymous):

thanks i got a couple more problems

OpenStudy (jojo4eva):

kk

OpenStudy (anonymous):

1/2 (12 dived by 3) + 3/5

OpenStudy (jojo4eva):

12/3=4 so 1/2(4)=2 so 2+3=5 vand 5/5=1 so you should get 1

OpenStudy (anonymous):

it has to be in fraction form a.57/8 b.19/2 c.17/2 d.22/3